# A deep learning model for chemical shieldings in molecular organic solids including anisotropy

### v2

@@ Abstract
Nuclear Magnetic Resonance (NMR) chemical shifts are powerful probes of local atomic and electronic structure that can be used to resolve the structures of powdered or amorphous molecular solids. Chemical shift driven structure elucidation depends critically on accurate and fast predictions of chemical shieldings, and machine learning (ML) models for shielding predictions are increasingly used as scalable and efficient surrogates for demanding ab initio calculations. However, the prediction accuracies of current ML models still lag behind those of the DFT reference methods they approximate. Here, we introduce ShiftML3, a deep-learning model that improves the accuracy of predictions of isotropic chemical shieldings in molecular solids, and does so while also predicting the full shielding tensor. On experimental benchmark sets, we find root-mean-squared errors with respect to experiment for ShiftML3 that approach those of DFT reference calculations, with RMSEs of 0.53 ppm for 1H, 2.4 ppm for 13C, and 7.2 ppm for 15N, compared to DFT values of 0.49 ppm, 2.3 ppm, and 5.8 ppm, respectively.
@@
In a nutshell, NMR crystallography consists in experimentally measuring and assigning chemical shifts for each of the atomic sites in a structure, and then comparing the measured shifts to those predicted by computation for a pool of candidate structures, that are themselves enumerated by a computational search. Chemical shifts are traditionally calculated using DFT methods. [7, 11, 12, 13, 14] This workflow has been successful for materials ranging from hybrid-perovskites [15, 16] and inorganic materials [17, 18, 19, 20, 21, 22] to enzyme active sites, [23, 24, 10] organic molecular solids, [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and notably for amorphous drug formulations. [40, 41, 42, 43] One of the key bottlenecks to address in this workflow is both the computational cost and the accuracy of chemical shift predictions.
@@
Here we use the training set from the most recent previous ShiftML model, and for which the computed shifts (computed at the PBE level of theory) contain the full tensor information. [49] This dataset is composed of molecular solids where the structures include (up to) 12 elements [H, C, N, O, F, Na, Mg, P, S, Cl, K, Ca] commonly found in organic compounds. It contains a total of 14,146 structures, including 7,548 sampled from the Cambridge structural database (CSD) via farthest point sampling (FPS) [70] to cover a broad range of structural motifs, and 6,598 additional thermally distorted structures. The ShiftML2 test set was divided into a new validation (505 structures, with 41,599 atomic environments) and a new test set (1,254 structures, with 106,230 atomic environments). In this split we chose a stratified selection strategy that ensures that both the distorted and the relaxed structure of any one CSD identifier are found either in the validation or the test set, avoiding data leakage between the sets. Furthermore, we ensure that both the new validation and new test sets contain at least a few environments of all 12 chemical elements. To compare the accuracies between ShiftML2 and ShiftML3 in a consistent fashion, the accuracies are reported on this new test set (referred to as the CSD-test set). More details of the training, validation and test sets for each element in CSD-test are given in section S1 of the supporting information. Due to computational constraints, ShiftML2 was trained on a representative subselection of atomic environments from the training set. The favorable computational scaling of neural network training via stochastic gradient descent allowed us to train ShiftML3 using all 1.4 million atomic environments contained in the 14,146 structures in the training set. The updated dataset containing the chemical shielding tensor data is available via a public data record associated with this publication. [71] We also use established datasets of experimental isotropic chemical shielding values, with 13 structures for 1H, 21 for 13C, and 15 for 15N, [49, 14] and a benchmark set of experimental 13C CSA principal values for 19 compounds. [72]
@@
In practice, nanoPET predicts the chemical shielding tensors in their irreducible spherical tensor (IST) representation in a manner similar to symmetry-adapted, rotationally-constrained models. [75] The chemical shielding tensors are decomposed into their irreducible real spherical representations: A scalar, $\lambda=0$ part, a $\lambda=1$ pseudovector, and a $\lambda=2$ proper 5-vector having the same rotational behaviour as their corresponding $Y^{\lambda}$ spherical harmonics. [76] We refer the reader to the following references, [77, 78, 79] for a more in depth discussion of the properties of the chemical shielding tensor and a discussion of its rotational properties. nanoPET uses a single message-passing sequence of unconstrained graph transformer modules, with one final linear readout layer per IST component, $T^{\lambda}$. The final tensor predictions are then obtained by a change of representation back from the predicted IST components to the tensor in the Cartesian basis.
During the training procedure, random rotations in their irreducible Wigner-D matrix representations are automatically applied to the target IST components at each training step, as well as random inversions. We train an ensemble consisting of 8 nanoPET models. The models are trained for 1,500 epochs, with a constant learning rate of $3\cdot 10^{-4}$. We train the model using mini batch gradient descent with the Adam optimizer [80] and a batch size of 4. The construction of an ensemble model has several benefits; in addition to increased prediction accuracies, an ensemble also provides uncertainty estimates and further reduces the rotational symmetry breaking of the nanoPET predictions. [81, 82] In Section S2 of the supporting information we make a detailed assessment of the residual symmetry errors of the ShiftML3 model.
@@
PET and most other atomistic neural networks make only point estimates of the target quantity. Constructing an ensemble is a universal strategy to equip ML models with uncertainty estimates, [84, 85] even if the model does not allow for a probabilistic interpretation of its predictions. An ensemble model is constructed by training multiple models. The final prediction for an atomic local environment, ${A}_{i}$, of an ensemble of Nens models, is then taken to be the mean of the individual predictions of the “committee” members, $\bar{y}(A_{i})$, and their spread is taken to be an indicator of the reliability of the model predictions:
@@
Figure 1: Parity plots between DFT-computed shieldings and predictions using ShiftML2 or ShiftML3 for (a,b) 1H, (c,d) 13C and (e,f) 15N evaluated on the new CSD-test set. The black line shows perfect correlation. Each point is colored according to the rescaled prediction uncertainty associated in each atomic site.
To evaluate the performance of ShiftML3, we first compare predicted isotropic chemical shieldings against DFT calculated isotropic shieldings on the CSD-test set. Figure 1 shows parity plots illustrating the overall agreement between ShiftML3 and DFT for the nuclei that are most often used in NMR crystallography. We observe an RMSE of 0.43 ppm, 2.3 ppm and 10.4 ppm for 1H, 13C and 15N, respectively. These correspond to an error reduction relative to ShiftML2 between 20% for 1H, to a factor of 2 for 13C. Similar levels of improvement are seen also for almost all nuclei represented in the dataset, as shown in Table 1 (with the exception of 39K, for which there are only 6 atomic enviroments in the test set, making the assessment inconclusive).
@@
We then evaluate the chemical shielding tensor predictions of the ShiftML3 model on the CSD-test set, and report the prediction accuracies of symmetric tensor components in Table 2. More information, including parity plots for 13C and 15N anisotropic shieldings can be found in section S5 of the supporting information. We report prediction accuracies for the combined principal components $\sigma_{\text{PAS}}$ and combined tensor components $\sigma_{ij}$. We report averaged prediction errors over all tensor components or principal components and across atomic environments in the test dataset, consistent with the error metric reported for chemical shieldings, which themselves are the average of the three principal components of the chemical shielding tensor. We obtain averaged eigenvalues of the ensemble, by first computing the eigendecomposition of the individual committee member predictions and then averaging over the committee prediction of principal components.
@@
The improvements in prediction accuracy against the DFT reference also translate to an increased accuracy of ShiftML predictions against experiments. We find that on a representative benchmark set of experimentally measured 1H and 13C shieldings, [49, 14] ShiftML3 predicts chemical shieldings with accuracies almost on par with GIPAW DFT, namely: 0.53(0.49) ppm for 1H and 2.4(2.3) ppm for 13C, for ShiftML3(DFT). Given that the errors of ShiftML3 against DFT are of similar size, this hints at substantial error cancellation. We also note that both computational predictions are affected by non-negligible error contributions due to the use of a static structure, neglecting finite-temperature effects and quantum fluctuations [90]. Figure 2, shows parity plots of isotropic chemical shieldings estimated with ShiftML2, ShiftML3 and DFT with respect to experimental chemical shifts. DFT values were evaluated recomputing the NMR shieldings using GIPAW [7] on the the Quantum ESPRESSO software [91, 92] with the same DFT parameters as in Ref. 49. Quantum Espresso GIPAW calculations for these 48 structures took approximately 80 hours on 32 cores of a modern compute node for all three benchmark sets. In contrast, ShiftML3 yielded shielding predictions for all structures in just 41 seconds on a personal computer (MacBook Pro M1), demonstrating speed-ups of over three orders of magnitude. The conversion between experimental chemical shifts and predicted shieldings is performed by fitting globally the shieldings across all test compounds to a linear form, also shown in figure 2. Further details about the experimental benchmark are discussed in section S7 of the supporting information.
We also assess the accuracy of the CSA predictions against experiments, based on a benchmark of 13C shieldings for a set of structures whose experimental shifts had been reported in Ref. 72. The accuracy in terms of the three principal components ($\sigma_{11},\sigma_{22},\sigma_{33}$) of the shielding tensor is on par with that of DFT. The RMSE of for ShiftML3 (5.2 ppm) is even smaller than for DFT (5.8 ppm) which is likely fortuitous (Fig. 3).
@@
To summarize our findings, we have presented a machine learning model for the prediction of chemical shieldings in organic crystals, using an unconstrained graph neural network architecture that delivers chemical shielding tensors and uncertainty quantification through an ensemble structure. We have shown that across all nuclei that are well-represented in the dataset, ShiftML3 predicts accurate shielding tensors and isotropic shieldings in organic crystals. The isotropic prediction RMSE of 0.43 ppm for 1H and 2.3 ppm for 13C against GIPAW reference calculations are a major improvement over previous ShiftML versions based on symmetry-adapted kernel models. The improved prediction accuracies translate to an accuracy against experimental benchmarks that practically matches that of DFT calculations, improving the confidence of NMR-guided structural determination. The availability of chemical shieldings anisotropy, with an accuracy that also approaches that of DFT, facilitates comparison with other types of NMR experiments, providing structural and mechanistic insights. The easy-to-use, fast, and openly-available implementation will facilitate usage for a broad community interested in the theoretical and experimental study of organic materials by means of NMR.
@@ Acknowledgements.
We thank Filippo Bigi and Michelangelo Domina for assistance with their nanoPET-IST implementation and Sergey Pozdnyakov for adapting the original PET architecture for isotropic shieldings. We thank Joseph William Abbott for the provision of a $\lambda$-SOAP model, which we used to protoype chemical shielding anisotropy predictions. This work has been supported by Swiss National Science Foundation Grant No. 200020_212046, and by the NCCR MARVEL, funded by the Swiss National Science Foundation (grant number 182892). MC also acknowledged funding from the ERC Horizon 2020 Grant No. 101001890-FIAMMA.
@@
The supporting information is provided below, containing additional details on dataset and model construction, uncertainty quantification, and the usage of the ShiftML3 Python package, which is available under an open source license at https://github.com/lab-cosmo/shiftml. The training, validation and testing datasets containing chemical shielding tensor data as well as training input files, model files and evaluation scripts will be made available upon publication via the public data record associated with this publication. [71]
@@
We evaluate the accuracy of ShiftML3 on the CSD-test set described in the main text. In Table S1, we report the size of the training, validation and CSD-test set together with the mean average error (MAE), root-mean-squared-error (RMSE), the coefficent of determination (R2 ), and the standardized RMSE ( % ) evaluation metrics of $\sigma_{\text{iso}}$ predictions of ShiftML3 for all chemical species contained in CSD-test.
@@
PET models learn the correct rotational behaviour of the targets via rotational augmentations. Langer and coworkers have investigated the effect relaxing exact rotational invariance to learned invariances, [74] and found that for realistic atomistic modelling tasks, this constraint can be safely relaxed. The remaining rotational fluctuations of the PET model and PET ensemble should however be carefully assessed for chemical shielding prediction tasks. Unlike the prediction of potential energies, or other structure wise global quantities, chemical shieldings are atom-wise properties. The common ansatz of predicting potential energies with neural networks, is by summing over local, per-atom terms. Shortcomings (imperfectly learned rotational invariance) of a non-invariant architecture might simply be compensated by error cancellation summing over atomic contributions. Since chemical shieldings and shielding tensors are exactly these local contribution (one prediction per atom), we study the effects of applying random rotations to structures and measuring the fluctuations of the predicted isotropic chemical shieldings. We apply 30 random rotations to each structure in the CSD test set and evaluate the standard deviation of the isotropic chemical shielding predictions. For a perfectly invariant model, these fluctuations should be exactly zero due to the invariance of the model. For an approximately invariant model, these predictions should be quite small, at most a fraction of the prediction RMSEs, such that they do not interfere with NMR crystal structure prediction workflows. In Table S2 we list rotational fluctuations, of ShiftML3 predictions across various chemical species and compare them with the rotational fluctuations of single PET models, highlighting the strongly reduced fluctuation of the ensemble of PET models that make ShiftML3.
@@
[…]
We evaluate the structure wise fluctuations of the RMSEs of the ShiftML3 predictions for CSD-test with respect to the GIPAW reference. To assess the effect of approximate invariance we also compute the fluctuations of the RMSE between all 1H environments in a structure and it non rotated structure. The mean fluctuation (mean of the standard deviation of fluctuations) of this RMSE across the CSD test set is 0.0065 ppm and the maximal fluctuation we find to be 0.0271 ppm.
@@
We benchmark the evaluation time of the ShiftML3 model. All benchmark timings are obtained via the ASE interface of the ShiftML library. We report timings that include setting up the ASE calculator once and then computing chemical shielding tensors for a series of structures. We believe that this protocol reflects most accurately the evaluation timings for a common NMR crystallography workflow. The timings via the ASE interface include also the neighbour list construction. The ASE interface does not support batched evaluations, or efficient parallelization over ensemble models for smaller system sizes. We report CPU-only and CPU+GPU timings. We report evaluation times on a Lenovo workstation with a AMD Ryzen Threadripper PRO 3955WX CPU with 16 cores. CPU+GPU timings are obtained on the same workstation, using a NVIDIA GeForce RTX 4070 Ti-super consumer GPU, which is representative of a configuration that does not require access to high-end AI-dedicated hardware. As shown in Table S3, even a low-end GPU accelerates by 4x-6x the evaluation of structures with the size of typical small-molecule pharmaceuticals. The acceleration is also evident at larger system sizes: the scaling of the ShiftML3 model for diamond supercells of increasing size is almost perfectly linear, and there is an order-of-magnitude acceleration when using a GPU (Fig. S2). These timings are also representative of large biomolecules, as hinted at by the inference timings for the protein crystal of 1PGA [93], the B1 immunoglobulin-binding domain of the streptococcal protein G with 56 amino acid residues and 3,656 atoms in the conventional cell (also shown in Fig. S2). 1PGA was a target of several solid-state NMR studies [94, 95] and therefore lies well within the range of atomistic systems of interest for which chemical shielding predictions may complement experimental studies. A computation of chemical shieldings for this protein takes 5.6 s using the GPU. Moving to even larger systems, we find that on the consumer GPU used for benchmarking with 16GB of VRAM, we can predict 13C shieldings in diamond supercells of up to 16,000 atoms before exhausting the graphics card video memory, demonstrating the memory efficiency of ShiftML3.
@@
To simplify usage of ShiftML3 for inference, we make it available with an easy-to-install Python package, released together with this publication. ShiftML is available from the Python Package Index PyPi and can be easily installed with

### v3

@@ Abstract
Abstract: Nuclear Magnetic Resonance (NMR) chemical shifts are powerful probes of local atomic and electronic structure that can be used to resolve the structures of powdered or amorphous molecular solids. Chemical shift driven structure elucidation depends critically on accurate and fast predictions of chemical shieldings, and machine learning (ML) models for shielding predictions are increasingly used as scalable and efficient surrogates for demanding ab initio calculations. However, the prediction accuracies of current ML models still lag behind those of the DFT reference methods they approximate. Here, we introduce ShiftML3, a deep-learning model that improves the accuracy of predictions of isotropic chemical shieldings in molecular solids, and does so while also predicting the full shielding tensor. On experimental benchmark sets, we find root-mean-squared errors with respect to experiment for ShiftML3 that approach those of DFT reference calculations, with RMSEs of 0.53 ppm for 1H, 2.4 ppm for 13C, and 7.2 ppm for 15N, compared to DFT values of 0.49 ppm, 2.3 ppm, and 5.8 ppm, respectively.
@@
In a nutshell, NMR crystallography consists of experimentally measuring and assigning chemical shifts for each of the atomic sites in a structure and then comparing the measured shifts to those predicted by computation for a pool of candidate structures, that are themselves enumerated by a computational search. Chemical shifts are traditionally calculated using DFT methods. [7, 11, 12, 13, 14] This workflow has been successful for materials ranging from hybrid-perovskites [15, 16] and inorganic materials [17, 18, 19, 20, 21, 22] to enzyme active sites, [23, 24, 10] organic molecular solids, [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and notably for amorphous drug formulations. [40, 41, 42, 43] One of the key bottlenecks to address in this workflow is both the computational cost and the accuracy of chemical shift predictions.
@@
Here we use the training set from the most recent previous ShiftML model, and for which the computed shifts (computed at the PBE level of theory) contain the full tensor information. [49] This dataset is composed of molecular solids where the structures include (up to) 12 elements [H, C, N, O, F, Na, Mg, P, S, Cl, K, Ca] commonly found in organic compounds. It contains a total of 14,146 structures, including 7,548 sampled from the Cambridge Structural Database (CSD) via farthest point sampling (FPS) [70] to cover a broad range of structural motifs, and 6,598 additional thermally distorted structures. The ShiftML2 test set was divided into a new validation (505 structures, with 41,599 atomic environments) and a new test set (1,254 structures, with 106,230 atomic environments). In this split we chose a stratified selection strategy that ensures that both the distorted and the relaxed structure of any one CSD identifier are found either in the validation or the test set, avoiding data leakage between the sets. Furthermore, we ensure that both the new validation and new test sets contain at least a few environments of all 12 chemical elements. To compare the accuracies between ShiftML2 and ShiftML3 in a consistent fashion, the accuracies are reported on this new test set (referred to as the CSD-test set). More details of the training, validation and test sets for each element in CSD-test are given in section S1 of the supporting information. Due to computational constraints, ShiftML2 was trained on a representative sub-selection of atomic environments from the training set. The favorable computational scaling of neural network training via stochastic gradient descent allowed us to train ShiftML3 using all 1.4 million atomic environments contained in the 14,146 structures in the training set. The updated dataset containing the chemical shielding tensor data is available via a public data record associated with this publication. [71] We also use established datasets of experimental isotropic chemical shielding values, with 13 structures for 1H, 21 for 13C, and 15 for 15N, [49, 14] and a benchmark set of experimental 13C CSA principal values for 19 compounds. [72]
@@
In practice, nanoPET predicts the chemical shielding tensors in their irreducible spherical tensor (IST) representation in a manner similar to symmetry-adapted, rotationally-constrained models. [75] The chemical shielding tensors are decomposed into their irreducible real spherical representations: A scalar, $\lambda=0$ part, a $\lambda=1$ pseudovector, and a $\lambda=2$ proper 5-vector having the same rotational behaviour as their corresponding $Y^{\lambda}$ spherical harmonics. [76] We refer the reader to the following references [77, 78, 79] for a more in-depth discussion of the properties of the chemical shielding tensor and a discussion of its rotational properties. nanoPET uses a single message-passing sequence of unconstrained graph transformer modules, with one final linear readout layer per IST component, $T^{\lambda}$. The final tensor predictions are then obtained by a change of representation back from the predicted IST components to the tensor in the Cartesian basis.
During the training procedure, random rotations in their irreducible Wigner-D matrix representations are automatically applied to the target IST components at each training step, as well as random inversions. We train an ensemble consisting of 8 nanoPET models. The models are trained for 1,500 epochs, with a constant learning rate of $3\cdot 10^{-4}$. We train the model using mini-batch gradient descent with the Adam optimizer [80] and a batch size of 4. The construction of an ensemble model has several benefits; in addition to increased prediction accuracies, an ensemble also provides uncertainty estimates and further reduces the rotational symmetry breaking of the nanoPET predictions. [81, 82] In Section S2 of the supporting information we make a detailed assessment of the residual symmetry errors of the ShiftML3 model.
@@
PET and most other atomistic neural networks make only point estimates of the target quantity. Constructing an ensemble is a universal strategy to equip ML models with uncertainty estimates, [84, 85] even if the model does not allow for a probabilistic interpretation of its predictions. An ensemble model is constructed by training multiple models. The final prediction for an atomic local environment, ${A}_{i}$, of an ensemble of Nens models, is then taken to be the mean $\bar{y}(A_{i})$ of the individual predictions $y^{k}(A_{i})$ of the “committee” members, and their spread is taken to be an indicator of the reliability of the model predictions:
@@
Figure 1: Parity plots between DFT-computed shieldings and predictions using ShiftML2 or ShiftML3 for (a,b) 1H, (c,d) 13C and (e,f) 15N evaluated on the new CSD-test set. The black line shows perfect correlation. Each point is colored according to the rescaled prediction uncertainty associated with each atomic site.
To evaluate the performance of ShiftML3, we first compare predicted isotropic chemical shieldings against DFT calculated isotropic shieldings on the CSD-test set. Figure 1 shows parity plots illustrating the overall agreement between ShiftML3 and DFT for the nuclei that are most often used in NMR crystallography. We observe an RMSE of 0.43 ppm, 2.3 ppm and 10.4 ppm for 1H, 13C and 15N, respectively. These correspond to an error reduction relative to ShiftML2 between 20% for 1H, to a factor of 2 for 13C. Similar levels of improvement are seen also for almost all nuclei represented in the dataset, as shown in Table 1 (with the exception of 39K, for which there are only 6 atomic environments in the test set, making the assessment inconclusive).
@@
We then evaluate the chemical shielding tensor predictions of the ShiftML3 model on the CSD-test set, and report the prediction accuracies of symmetric tensor components in Table 2. More information, including parity plots for 13C and 15N anisotropic shieldings can be found in section S5 of the supporting information. We report prediction accuracies for the combined principal components $\sigma_{\text{PAS}}$ (principal axis system) and combined tensor components $\sigma_{ij}$. We report averaged prediction errors over all tensor components or principal components and across atomic environments in the test dataset, consistent with the error metric reported for chemical shieldings, which themselves are the average of the three principal components of the chemical shielding tensor. We obtain averaged eigenvalues of the ensemble by first computing the eigen decomposition of the individual committee member predictions and then averaging over the committee predictions of principal components.
@@
The improvements in prediction accuracy against the DFT reference also translate to an increased accuracy of ShiftML predictions against experiments. We find that on a representative benchmark set of experimentally measured 1H and 13C shieldings, [49, 14] ShiftML3 predicts chemical shieldings with accuracies almost on par with GIPAW DFT, namely: 0.53(0.49) ppm for 1H and 2.4(2.3) ppm for 13C, for ShiftML3(DFT). Given that the errors of ShiftML3 against DFT are of similar size, this hints at substantial error cancellation. We also note that both computational predictions are affected by non-negligible error contributions due to the use of a static structure, neglecting finite-temperature effects and quantum fluctuations [90]. Figure 2 shows parity plots of isotropic chemical shieldings estimated with ShiftML2, ShiftML3 and DFT with respect to experimental chemical shifts. DFT values were evaluated by recomputing the NMR shieldings using GIPAW [7] on the Quantum ESPRESSO software [91, 92] with the same DFT parameters as in Ref. 49. Quantum ESPRESSO GIPAW calculations for these 48 structures took approximately 80 hours on 32 cores of a modern compute node for all three benchmark sets. In contrast, ShiftML3 yielded shielding predictions for all structures in just 41 seconds on a personal computer (MacBook Pro M1), demonstrating speed-ups of over three orders of magnitude. The conversion between experimental chemical shifts and predicted shieldings is performed by fitting globally the shieldings across all test compounds to a linear form, also shown in figure 2. Further details about the experimental benchmark are discussed in section S7 of the supporting information.
We also assess the accuracy of the CSA predictions against experiments, based on a benchmark of 13C shieldings for a set of structures whose experimental shifts had been reported in Ref. 72. The accuracy in terms of the three principal components ($\sigma_{11},\sigma_{22},\sigma_{33}$) of the shielding tensor is on par with that of DFT. The RMSE for ShiftML3 (5.2 ppm) is even smaller than for DFT (5.8 ppm) which is likely fortuitous (Fig. 3).
@@
To summarize our findings, we have presented a machine learning model for the prediction of chemical shieldings in organic crystals, using an unconstrained graph neural network architecture that delivers chemical shielding tensors and uncertainty quantification through an ensemble structure. We have shown that across all nuclei that are well-represented in the dataset, ShiftML3 predicts accurate shielding tensors and isotropic shieldings in organic crystals. The isotropic prediction RMSE of 0.43 ppm for 1H and 2.3 ppm for 13C against GIPAW reference calculations are a major improvement over previous ShiftML versions based on symmetry-adapted kernel models. The improved prediction accuracies translate to an accuracy against experimental benchmarks that practically matches that of DFT calculations, improving the confidence of NMR-guided structural determination. The availability of chemical shielding anisotropy, with an accuracy that also approaches that of DFT, facilitates comparison with other types of NMR experiments, providing structural and mechanistic insights. The easy-to-use, fast, and openly available implementation will facilitate usage for a broad community interested in the theoretical and experimental study of organic materials by means of NMR.
@@ Acknowledgements.
We thank Filippo Bigi and Michelangelo Domina for assistance with their nanoPET-IST implementation and Sergey Pozdnyakov for adapting the original PET architecture for isotropic shieldings. We thank Joseph William Abbott for the provision of a $\lambda$-SOAP model, which we used to prototype chemical shielding anisotropy predictions. This work has been supported by Swiss National Science Foundation Grant No. 200020_212046, and by the NCCR MARVEL, funded by the Swiss National Science Foundation (grant number 182892). MC also acknowledged funding from the ERC Horizon 2020 Grant No. 101001890-FIAMMA.
@@
The supporting information is provided below, containing additional details on dataset and model construction, uncertainty quantification, and the usage of the ShiftML3 Python package, which is available under an open-source license at https://github.com/lab-cosmo/shiftml. The training, validation and testing datasets containing chemical shielding tensor data as well as training input files, model files and evaluation scripts will be made available upon publication via the public data record associated with this publication. [71]
@@
We evaluate the accuracy of ShiftML3 on the CSD-test set described in the main text. In Table S1, we report the size of the training, validation and CSD-test set together with the mean absolute error (MAE), root-mean-squared-error (RMSE), the coefficient of determination (R2), and the standardized RMSE (%) evaluation metrics of $\sigma_{\text{iso}}$ predictions of ShiftML3 for all chemical species contained in CSD-test.
@@
PET models learn the correct rotational behaviour of the targets via rotational augmentations. Langer and coworkers have investigated the effect of relaxing exact rotational invariance to learned invariances, [74] and found that for realistic atomistic modelling tasks, this constraint can be safely relaxed. The remaining rotational fluctuations of the PET model and PET ensemble should however be carefully assessed for chemical shielding prediction tasks. Unlike the prediction of potential energies, or other structure-wise global quantities, chemical shieldings are atom-wise properties. The common ansatz of predicting potential energies with neural networks, is by summing over local, per-atom terms. Shortcomings (imperfectly learned rotational invariance) of a non-invariant architecture might simply be compensated by error cancellation summing over atomic contributions. Since chemical shieldings and shielding tensors are exactly these local contributions (one prediction per atom), we study the effects of applying random rotations to structures and measuring the fluctuations of the predicted isotropic chemical shieldings. We apply 30 random rotations to each structure in the CSD-test set and evaluate the standard deviation of the isotropic chemical shielding predictions. For a perfectly invariant model, these fluctuations should be exactly zero due to the invariance of the model. For an approximately invariant model, these predictions should be quite small, at most a fraction of the prediction RMSEs, such that they do not interfere with NMR crystal structure prediction workflows. In Table S2 we list rotational fluctuations of ShiftML3 predictions across various chemical species and compare them with the rotational fluctuations of single PET models, highlighting the strongly reduced fluctuation of the ensemble of PET models that make up ShiftML3.
@@
Figure S1: Correlations of uncertainty estimates $\text{std}_{\text{pred}}[\sigma_{\text{iso}}]$ across all atomic environments in the CSD-test set with the rotational fluctuations of the same environments std[$_{\text{rot}}[\sigma_{\text{iso}}$], for 1H (a), 13C (b), 15N (c) and 17O (d).
We evaluate the structure-wise fluctuations of the RMSEs of the ShiftML3 predictions for CSD-test with respect to the GIPAW reference. To assess the effect of approximate invariance we also compute the fluctuations of the RMSE between all 1H environments in a structure and its non-rotated structure. The mean fluctuation (mean of the standard deviation of fluctuations) of this RMSE across the CSD-test set is 0.0065 ppm and the maximal fluctuation we find to be 0.0271 ppm.
@@
We benchmark the evaluation time of the ShiftML3 model. All benchmark timings are obtained via the ASE interface of the ShiftML library. We report timings that include setting up the ASE calculator once and then computing chemical shielding tensors for a series of structures. We believe that this protocol reflects most accurately the evaluation timings for a common NMR crystallography workflow. The timings via the ASE interface include also the neighbour list construction. The ASE interface does not support batched evaluations, or efficient parallelization over ensemble models for smaller system sizes. We report CPU-only and CPU+GPU timings. We report evaluation times on a Lenovo workstation with an AMD Ryzen Threadripper PRO 3955WX CPU with 16 cores. CPU+GPU timings are obtained on the same workstation, using an NVIDIA GeForce RTX 4070 Ti SUPER consumer GPU, which is representative of a configuration that does not require access to high-end AI-dedicated hardware. As shown in Table S3, even a low-end GPU accelerates by 3x-4x the evaluation of structures with the size of typical small-molecule pharmaceuticals. The acceleration is also evident at larger system sizes: the scaling of the ShiftML3 model for diamond supercells of increasing size is almost perfectly linear, and there is an order-of-magnitude acceleration when using a GPU (Fig. S2). These timings are also representative of large biomolecules, as hinted at by the inference timings for the protein crystal of 1PGA [93], the B1 immunoglobulin-binding domain of the streptococcal protein G with 56 amino acid residues and 3,656 atoms in the conventional cell (also shown in Figure S2). 1PGA was a target of several solid-state NMR studies [94, 95] and therefore lies well within the range of atomistic systems of interest for which chemical shielding predictions may complement experimental studies. A computation of chemical shieldings for this protein takes 5.6 s using the GPU. Moving to even larger systems, we find that on the consumer GPU used for benchmarking with 16 GB of VRAM, we can predict 13C shieldings in diamond supercells of up to 16,000 atoms before exhausting the graphics card video memory, demonstrating the memory efficiency of ShiftML3.
@@
To simplify usage of ShiftML3 for inference, we make it available with an easy-to-install Python package, released together with this publication. ShiftML is available from the Python Package Index (PyPI) and can be easily installed with